Carl and Dennis are brothers. The sum of their age is 29. In 6 years time Carl will be the same age as Dennis was 3 years ago.
Determine the current age of the two brothers.

Answer by mananth(16946)

Carl and Dennis are brothers. The sum of their age is 29.
C+D = 29--------------------(1)
In 6 years time Carl will be the same age as Dennis was 3 years ago.
(c+6) = (D-3)
C-D=-9-----------------------(2)
Determine the current age of the two brothers.
Add 1 & 2
2C=20
C=10
C+D=29
D=19
current age of the two brothers Carl 10 years, Dennis 19 years